OpenStudy (anonymous):

The lengths of two sides of a right triangle are given. Find the length of the third side. Round to the nearest tenth if necessary. legs: 28 in. and 15 in. A. 37.5 in. B. 23.6 in. C. 29.6 in. D. 31.8 in.

OpenStudy (anonymous):

Use the Pythagorean theorem to find the remaining side: 28^2 + 15^2 = C^2 So: 1009 = C^2 Take the square root: C = 31.76476... So the answer is D.

OpenStudy (anonymous):

\[Hypotenuse^2 = Base^2 + Hieght^2\]